• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name B. Quinn

    Log entry time 16:09:28 on June 4,2010

    Entry number 324519

    keyword=Day shift summary

    08:00 Start of shift with experts finishing tests. Pb/diamond target #3 @50uA

    08:50 Luis found Lumis were off, turned on

    09:00 Tried to start a run using Qweak crate (which had been rebooted by Bob Michaels) no luck.

    09:40 No (or almost no) beam in last 25 minutes. Turned off cooler and turned on heater. (MCC predicts beam within a few minutes.)
    10:10 MCC was wrong... but they now ask if we are ready for beam back.
    10:57 did beam recovery procedure and started to ramp up current about 15 minutes ago. MCC having problems. They now advise they are having trouble with our fast feedback. Expect to have it fixed in a few minutes.

    12:15 RC requests that we get an escorted controlled access at 2pm to allow moller magnet to be re-filled. Also during the down, we should have MCC reboot IOC HA1 to see if it fixes problem with Harp 4B.
    In case calibration coefficients are lost during reboot, they are:
    IBC1H00: 52.950 IBC1H00A: 44.770

    13:30 RC calls to say Moller refill has been re-scheduled for 8pm. We'll continue production running.

    16:00 Learned that we don't have communication with Right arm HV (log entry 32513) because Luis had taken control during the 3 o'clock meeting. Still not resolved.... Luis should restore our control when he's done with it.

    Start Run OK
    time number Multiplets (k)
    09:04 4259 20.4
    09:55 4260 0 (no beam)
    10:51 4261 48.5
    11:42 4262 86.1
    12:34 4263 76.3
    13:25 4264 71.6
    14:16 4265 52.3
    15:59 4266